A researcher is developing a regression equation to predict income (Y) in thousands of dollars from age (X1) and educational status (X2). Educational status is a dummy variable with 0 = high school or less and 1 = some education beyond high school. Y = 11.4 + 0.6X1 + 4.7X2 + e. What is the predicted income for a 26-year-old with a college degree?

a.)$27,700

b.)$31,700

c.)$41,860

d.)$24,220
Answer #1

Ans:

Regression equation:

Y = 11.4 + 0.6X1 + 4.7X2 + e

When X1=26

X2=1

Then ,predicted salary

Y=11.4 + 0.6*26+ 4.7*1=31.7 thousand dollars

Y=31700 dollars

Option B is correct.